A steel rod is initially at 20°C and has a length of 2m. It has a cross-sectional area of 0.001m^2 , If it is heated to 120°C by how much does its length increase?


1
Expert's answer
2021-07-01T15:29:44-0400

"l=l_0(1+\alpha \Delta t)=2\cdot(1+2.4\cdot10^{-5}\cdot100)=2.0048\ (m)"


"\Delta l=0.0048\ (m)"